simplify if possible
32u^7z^6/8u^3z^4

Answer by jessica43(140)

To solve this problem, you need to use one of the basic rules of exponents: Whenever you divide two terms with the same base, you can subtract the exponents:
(x^m)/(x^n) = x^(m-n)
So for this problem, you can use this rule to simplify:
(32u^7z^6)/(8u^3z^4)
(32u^4z^6)/(8z^4) (used the rule: (u^7)/(u^3) = u^(7-3) = u^4))
(32u^4z^2)/(8) (used the rule: (z^6)/(z^4) = z^(6-4) = z^2)
4u^4z^2 (divided 32 by 8)
So your simplified answer is 4u^4z^2.